# Self-Harmonized Chain of Thought

### v1

@@ Abstract
Chain-of-thought (CoT) prompting reveals that large language models are capable of performing complex reasoning via intermediate steps. CoT prompting is mainly categorized into two paradigms. The first approach utilizes straightforward prompts like “Let’s think step by step” to generate a sequential thought process before yielding an answer. The second approach makes use of human-crafted, step-by-step demonstrations to guide the model’s reasoning process. The latter method replaces manually written demonstrations with ones generated by large language models using “Let’s think step by step”. This approach sometimes leads to reasoning errors, highlighting the need to diversify demonstrations to mitigate its misleading effects. However, diverse demonstrations pose challenges for effective representations. In this work, we propose ECHO, a self-harmonized chain-of-thought prompting method. Consolidating diverse solution paths into a uniform and effective solution pattern, ECHO demonstrates the best overall performance in three reasoning domains [^1] Our code is available: https://github.com/Xalp/ECHO.
@@ 1 Introduction
A recent technique known as chain-of-thought (CoT) prompting Wei et al. (2022) has enhanced the ability of large language models (LLMs) to tackle complex reasoning tasks. By decomposing complex problems into a series of intermediate steps, CoT prompting effectively guides LLMs in their reasoning process.
@@
CoT methods in large language models typically use two prompting paradigms: Zero-shot-CoT Kojima et al. (2022) and Few-shot-CoT Wei et al. (2022). Zero-shot-CoT employs a single, universal prompt like “Let’s think step by step” to guide models in forming reasoning chains without needing specific examples, enhancing their zero-shot reasoning capabilities. The second paradigm, Few-shot-CoT, involves few-shot prompting with human-crafted demonstrations, pairing each question with a detailed reasoning chain comprising multiple intermediate steps that lead to the final answer.
Few-Shot-CoT necessitates the creation of human-crafted examples for each specific domain, which can be time-consuming and expensive. To address this, a method called Auto-CoT Zhang et al. (2023) has been developed. Auto-CoT automates the process of creating demonstrations by initially employing Zero-shot-CoT. Specifically, it clusters questions based on their similarity and selects a representative question from each cluster. Then, using the prompt “Let’s think step by step,” Auto-CoT generates the rationale for each chosen question. This automated process efficiently forms demonstrations and mitigates the manual effort required in Few-shot-CoT. Although it offers a solution to the need for human-crafted examples of Few-Shot-CoT, Auto-CoT faces its own challenge known as misleading by similarity Zhang et al. (2023)as the method depends on demonstrations initialized by Zero-shot-CoT. Some of these demonstrations may contain incorrect reasoning processes and answers. If one of the wrong demonstrations happens to be similar to the problem to be solved, it can be misleading. To make the method more robust, Zhang et al. (2023) select one from each question cluster to maximize diversity. Although this diversity mitigates the misleading effect of similar demonstrations, it introduces two new problems. First, some demonstrations may be too dissimilar or irrelevant to the actual question, making them ineffective. Second, diverse demonstrations might encompass varied solution patterns, including nuanced ones, making them less representative to learn from.
To address the two problems brought about by diverse demonstrations, we propose a novel method that aims to unify diverse rationale patterns into one general pattern, which is the balance point between all patterns. We name our method ECHO (Self-Harmonized Chain of Thought). It consists of three main steps: First, split a given dataset into several clusters and select one representative question from each cluster. Second, use Zero-shot-CoT for each question to generate its rationale. Third, employ a dynamic prompting mechanism where demonstrations can be improved by each other. In this step, one demonstration is randomly chosen to regenerate its rationale at each iteration, while the remaining demonstrations serve as in-context examples. This process ensures continuous refinement and adaptation of reasoning patterns between demonstrations.
@@
We conducted experiments across three different reasoning domains. ECHO demonstrates better overall performance (+2.8%) than other baselines. We also conduct comprehensive ablation studies to understand why unifying diversity improves performance.
We summarize our main contributions in this work as follows:
@@
  • We propose ECHO, a novel approach that automatically improves the quality of demonstrations in the CoT process by unifying the diversity.
@@
  • We design an iterative unifying prompting framework that is generally effective across various tasks via reducing the demonstrations’ variety.
@@
  • Extensive experiments have shown that our proposed approach offers competitive results on arithmetic, commonsense and symbolic reasoning domains by reducing diversity.
@@ 2 Related Work
CoT prompting, a gradient-free method, effectively encourages LLMs to generate intermediate reasoning steps before yielding the final answer. Wei et al. (2022) proposes Few-shot-CoT, which utilizes effective manual demonstrations to facilitate CoT reasoning. Later, Kojima et al. (2022) improved this method for zero-shot reasoning.
Recent studies have focused on creating more complex demonstrations or employing ensemble-like strategies. For example, some prompting methods Zhou et al. (2022); Wang et al. (2023); Yao et al. (2023) adopt a problem decomposition approach, breaking down complex problems into simpler subproblems. Self-consistency and some other CoT methods Wang et al. (2022); Asai et al. (2023) involve reasoning from multiple paths. Moreover, some prompting methods can generate executable programs to aid in the computation process within the arithmetic domain Gao et al. (2022); Chen et al. (2022); Pi et al. (2022).
Although designing increasingly complex prompts improves CoT effectiveness, it also increases the human efforts involved. Surprisingly, little attention has been paid to automating prompt generation. Among the few efforts in this area, Auto-CoT Zhang et al. (2023) stands out. Auto-CoT automates the process by clustering a dataset and selecting a representative question from each cluster, followed by using Zero-shot-CoT to generate rationales for these questions.
A key consideration in Auto-CoT Zhang et al. (2023) is the avoidance of using questions in the same cluster as demonstrations. The rationale behind this choice is rooted in the observation that if demonstrations are too similar to the target question, there is a higher risk that the model might replicate its mistakes, since Zero-shot-CoT Kojima et al. (2022) does not guarantee correctness in its responses (i.e. the demonstrations might be wrong). However, the use of diverse demonstrations might include irrelevant demonstrations or non-representative solution patterns. Our approach is designed to address these potential limitations.
@@ 3 Self-Harmonized Chain of Thought
Our method involves three main steps as shown in Figure 2: (1) question clustering: partition questions of a given dataset into a few clusters based on their similarity; (2) demonstration sampling: select a representative question from each cluster and generate its reasoning chain using Zero-shot-CoT. These first two steps are similar to Auto-CoT with minor differences Zhang et al. (2023). (3) demonstration unification: one demonstration is randomly selected for rationale update in each iteration, while the remaining demonstrations serve as in-context examples. This unification process forces each rationale to learn from the remaining ones to build a coherent pattern. The process iteratively cycles through each demonstration once per iteration and continues over multiple iterations to achieve consistency across all demonstrations.
Question Clustering
In our method, we first apply clustering to a set of questions $Q$. Each question is transformed into a fixed-size vector representation using Sentence-BERT Reimers and Gurevych (2019). Following this, we employ a $k$-means clustering model to categorize these vector representations into $k$ distinct clusters. Within each cluster $i$, the questions are ordered in a list $\mathbf{q}^{(i)}=[q_{1}^{(i)},q_{2}^{(i)},\ldots]$, based on their distance to the centroid of the cluster. This methodology is comprehensively detailed in lines 1 through 5 of the algorithm. In the Auto-CoT framework, Zhang et al. (2023), the number of clusters is equal to the number of output demonstrations, and one representative of each cluster is chosen as the demonstration. However, our approach adopts a different strategy by allowing for a greater number of clusters. This expansion enables the inclusion of a wider range of demonstrations in the unification process. This strategy has two benefits: It not only facilitates effective learning from more diverse patterns with the additional demonstrations, but also ensures that the final pattern can be applied to a wider range of demonstrations.
Demonstration Sampling

### v2

@@ Abstract
Chain-of-thought (CoT) prompting has demonstrated the capacity of large language models to perform complex reasoning through intermediate steps. While effective, current CoT methods face challenges: Zero-shot-CoT can lead to reasoning errors, and Few-shot-CoT requires labor-intensive manual demonstrations. Auto-CoT attempts to address these issues by automatically generating diverse demonstrations, but this diversity can lead to inconsistent reasoning patterns. We propose ECHO (Self-Harmonized Chain of Thought), a novel method that unifies diverse solution paths into a consistent and effective reasoning pattern. ECHO employs an iterative process to refine and harmonize automatically generated demonstrations, mitigating the limitations of existing approaches. Our comprehensive experiments across arithmetic, commonsense, and symbolic reasoning tasks demonstrate that ECHO outperforms Auto-CoT by an average of 2.8%. These findings suggest that ECHO represents a significant step towards more robust and generalizable automated reasoning in large language models. [^1] Our code is available: https://github.com/Xalp/ECHO.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in various natural language processing tasks. However, their performance on complex reasoning tasks has been a persistent challenge. A recent technique known as chain-of-thought (CoT) prompting Wei et al. (2022) has significantly enhanced LLMs’ ability to tackle such tasks by decomposing complex problems into a series of intermediate steps.
@@
CoT methods typically employ two prompting paradigms: Zero-shot-CoT Kojima et al. (2022) and Few-shot-CoT Wei et al. (2022). Zero-shot-CoT uses a universal prompt like “Let’s think step by step” to guide models in forming reasoning chains without specific examples. Few-shot-CoT, on the other hand, involves few-shot prompting with human-crafted demonstrations, pairing each question with a detailed reasoning chain.
While Few-shot-CoT has shown promising results, it requires the creation of human-crafted examples for each specific domain, which can be time-consuming and expensive. To address this limitation, Auto-CoT Zhang et al. (2023) was developed to automate the process of creating demonstrations by employing Zero-shot-CoT. However, Auto-CoT faces its own challenges:
  • Misleading by similarity: Some demonstrations generated by Zero-shot-CoT may contain incorrect reasoning processes or answers, potentially misleading the model when solving similar problems.
  • Ineffective diversity: To mitigate the misleading effect, Auto-CoT selects diverse demonstrations. However, this approach can lead to demonstrations that are too dissimilar or irrelevant to the actual question, reducing their effectiveness.
  • Inconsistent solution patterns: Diverse demonstrations might encompass varied solution patterns, including nuanced ones, making them less representative for the model to learn from.
To address these limitations, we propose ECHO (sElf-Harmonized Chain of Thought), a novel method that aims to unify diverse rationale patterns into one general pattern. Our approach is inspired by Cognitive Load Theory Sweller (1988), which posits that learning is most effective when the cognitive load on working memory is minimized. By unifying demonstrations, ECHO creates a more coherent set of examples, potentially reducing the cognitive load, thus facilitating more effective reasoning patterns that can be easily followed.
ECHO consists of three main steps: (1) Cluster a given dataset and select representative questions. (2) Generate initial rationales using Zero-shot-CoT. (3) Employ a dynamic prompting mechanism to iteratively improve demonstrations using each other as in-context examples.
This iterative process not only refines the quality of individual demonstrations but also promotes consistency across the entire set, potentially reducing the cognitive load required to process and apply the demonstrated reasoning patterns.
@@
Our experiments across three different reasoning domains demonstrate that ECHO outperforms other baselines by 2.8% overall. We also conduct comprehensive ablation studies to understand the benefits of unifying diversity in improving performance. These results support the hypothesis that a more unified set of demonstrations, as suggested by Cognitive Load Theory, can indeed lead to improved reasoning capabilities in LLMs.
The main contributions of this work are:
@@
  • A novel approach (ECHO) that automatically improves the quality of demonstrations in the CoT process by unifying diversity.
@@
  • An iterative unifying prompting framework that is effective across various tasks by reducing demonstration variety.
@@
  • Extensive experiments showing competitive results on arithmetic, commonsense, and symbolic reasoning domains through diversity reduction, supporting the efficacy of our cognitively-informed approach.
@@ 2 Related Work
Chain-of-Thought (CoT) Prompting
CoT prompting has emerged as a powerful, gradient-free method for enhancing the reasoning capabilities of Large Language Models (LLMs). Wei et al. (2022) introduced Few-shot-CoT, which utilizes manually crafted demonstrations to guide the model’s reasoning process. Kojima et al. (2022) then proposed Zero-shot-CoT, extending the method to zero-shot scenarios.
Recent studies have focused on creating more complex demonstrations or employing ensemble-like strategies. For example, some prompting methods Zhou et al. (2022); Wang et al. (2023); Yao et al. (2023) adopt a problem decomposition approach, breaking down complex problems into simpler subproblems. Self-consistency and other CoT methods Wang et al. (2022); Asai et al. (2023); Sun et al. involve reasoning from multiple paths or iteratively. Moreover, some prompting methods can generate executable programs to aid in the computation process within the arithmetic domain Gao et al. (2022); Chen et al. (2022); Pi et al. (2022).
Automated Prompt Generation
Although designing increasingly complex prompts improves CoT effectiveness, it also increases the human effort involved. Auto-CoT Zhang et al. (2023) automates the process by clustering a dataset and selecting a representative question from each cluster, followed by using Zero-shot-CoT to generate rationales for these questions.
A key consideration in Auto-CoT Zhang et al. (2023) is the avoidance of using questions in the same cluster as demonstrations. The rationale behind this choice is rooted in the observation that if demonstrations are too similar to the target question, there is a higher risk that the model might replicate its mistakes, since Zero-shot-CoT Kojima et al. (2022) does not guarantee correctness in its responses (i.e., the demonstrations might be wrong). However, the use of diverse demonstrations might include irrelevant demonstrations or non-representative solution patterns.
@@ 3 Self-Harmonized Chain of Thought
Our method, ECHO, involves three main steps as shown in Figure 2:
  1. 1.
    Question clustering: Partition questions of a given dataset into clusters based on their similarity.
  2. 2.
    Demonstration sampling: Select a representative question from each cluster and generate its reasoning chain using Zero-shot-CoT.
  3. 3.
    Demonstration unification: Iteratively update rationales to build a coherent pattern across all demonstrations.
These steps are designed to address the limitations of previous methods by creating a unified, representative set of demonstrations. The first two steps are similar to Auto-CoT Zhang et al. (2023) with minor differences, while the third step is unique to our approach.
3.1 Question Clustering
We begin by applying clustering to a set of questions $Q$. Each question is transformed into a fixed-size vector representation using Sentence-BERT Reimers and Gurevych (2019), which provides semantically meaningful embeddings. We then employ a $k$-means clustering model to categorize these vector representations into $k$ distinct clusters. Within each cluster $i$, the questions are ordered in a list $\mathbf{q}^{(i)}=[q_{1}^{(i)},q_{2}^{(i)},\ldots]$, based on their distance to the centroid of the cluster.
Unlike Auto-CoT Zhang et al. (2023), where the number of clusters equals the number of output demonstrations, our approach allows for a greater number of clusters. This expansion enables the inclusion of a wider range of demonstrations in the unification process, facilitating effective learning from more diverse patterns and ensuring that the final pattern can be applied to a wider range of demonstrations.
3.2 Demonstration Sampling